# Beyond the Kolmogorov Barrier: A Learnable Weighted Hybrid Autoencoder for Model Order Reduction

### v3

@@ Abstract
Representation learning for high-dimensional, complex physical systems aims to identify a low-dimensional intrinsic latent space, which is crucial for reduced-order modeling and modal analysis. To overcome the well-known Kolmogorov barrier, deep autoencoders (AEs) have been introduced in recent years, but they often suffer from poor convergence behavior as the rank of the latent space increases. To address this issue, we propose the learnable weighted hybrid autoencoder, a hybrid approach that combines the strengths of singular value decomposition (SVD) with deep autoencoders through a learnable weighted framework. We find that the introduction of learnable weighting parameters is essential — without them, the resulting model would either collapse into a standard POD or fail to exhibit the desired convergence behavior. Interestingly, we empirically find that our trained model has a sharpness thousands of times smaller compared to other models. Our experiments on classical chaotic PDE systems, including the 1D Kuramoto-Sivashinsky and forced isotropic turbulence datasets, demonstrate that our approach significantly improves generalization performance compared to several competing methods. Additionally, when combining with time series modeling techniques (e.g., Koopman operator, LSTM), the proposed technique offers significant improvements for surrogate modeling of high-dimensional multi-scale PDE systems.
keywords:
Reduced-order modeling , Dimensionality Reduction , POD , Deep Learning, Surrogate Modeling
[…]
[1]organization=Department of Mechanical, Aerospace, and Nuclear Engineering, Rensselaer Polytechnic Institute, addressline=110 8th St, city=Troy, postcode=12180, state=New York, country=USA
[…]
[2]organization=Scientific Computation Research Center, Rensselaer Polytechnic Institute, addressline=110 8th St, city=Troy, postcode=12180, state=New York, country=USA
@@ 1 Introduction
Computational fluid dynamics involves solving large dynamical systems with millions of degrees of freedom, resulting in significant computational overhead. In order to alleviate this problem, reduced-order modeling [1] is widely used, which uses a smaller number of modes to provide an approximate solution at a lower computational expense. A crucial step in reduced-order modeling is the projection of system states to a reduced latent space. The projection error on the latent space represents the overhead error that affects the solution accuracy throughout the time-stepping scheme. Linear dimensionality reduction techniques such as Proper Orthogonal Decomposition (POD) [2, 3, 4] are often used to create efficient representations of large-scale systems by projecting the solution manifold onto the space spanned by a set of linear orthonormal basis. Advances in deep learning techniques, such as deep autoencoders (AE) [5], capture intrinsic non-linear features for better compression and retrieval of high-fidelity information, outperforming POD in lower ranks, thus overcoming the well-known Kolmogorov barrier [6, 7]. Milano and Koumoutsakos [8] highlighted one of the first works to utilize a fully connected autoencoder to reconstruct the flow field, offering better performance compared to POD. Further studies have reported the usage of convolutional neural networks on 2D or 3D flow fields [9, 10, 11]. Such methods have been adopted in the fluid dynamics community to obtain a nonlinear model order reduction [12], but they do not provide rapid error convergence at higher rank. Recently, there have also been studies on using the hybrid approach [13, 14] combining POD with deep learning, by passing the latent space produced by POD to a neural network to find the corrections required to enhance reconstruction. These hybrid techniques have proven to enhance reconstruction beyond the capabilities of vanilla autoencoders. Following a similar path, here we propose a novel dimensionality reduction technique that combines traditional dimensionality reduction technique, i.e., POD, with deep learning techniques in a weighted manner at the encoder and decoder stage, where such weights of hybridization are also learned from data, to achieve a more effective and flexible dimensionality reduction. We also compare this approach with a straightforward hybrid technique, where a direct sum of POD and AE is utilized to construct the autoencoder, demonstrating the need of using learnable weighting parameters between POD and AE. Interestingly, our proposed approach obtains flat minima as opposed to other approaches, which contributes to the improved generalization and noise robustness.
@@
Building on our proposed framework, we also demonstrate the application of our framework to surrogate modeling tasks, particularly in the context of system evolution. We explore two distinct applications of our framework. First, we integrate our framework with the Koopman operator, where our framework facilitates the mapping of Koopman embedding back to physical space, while learning a linear forward model to evolve the system states in a reduced space. Second, we leverage long-short-term memory (LSTM) for system evolution [15]. LSTMs are usually trained on reduced states, obtained via POD [16, 17] or a non-linear autoencoder [18, 19] to serve as a surrogate for the temporal evolution of latent states. The computational costs can be prohibitive when using LSTMs on full order states, which can typically be in the range of millions for complex scenarios in fluid dynamics. By integrating our improved dimensionality reduction technique with a forward model for the system (linear or non-linear), we show that the enhanced latent space leads to more accurate prediction of the system dynamics and provide insights into the individual error contributions from the dimensionality reduction and time series model. This underscores the fact that the quality of the reduced representation is the primary bottleneck in achieving highly accurate predictions, as suboptimal latent spaces inherently limit the effectiveness of any downstream models.
@@
The remainder of the manuscript is structured as follows. In Section 2 we present the details of our hybrid dimensionality reduction technique, Koopman operator used for long term forecast and nonlinear surrogate for learning PDE dynamics. Section 3 describes the datasets used in each of the tasks, network architecture, and training hyper-parameters. Section 4 provides a comprehensive evaluation of the proposed framework with respect to pure dimensionality reduction tasks and other downstream applications, comparing its performance against other techniques, and examining some key properties related to sharpness and noise robustness. Finally, Section 5 concludes the paper.
@@ 2.1 Dimensionality Reduction
    the neural network encoder with output dimension as $r$, denoted as $\phi_{\textrm{NN}}$. As shown in eq. 1, the latent state $\mathbf{z}\in\mathbb{R}^{r}$ is obtained by a weighted sum of POD projection and the output of encoder,
    $\mathbf{z}=(\mathbf{1}-\mathbf{a})\phi_{\textrm{POD}}(\mathbf{x})+\mathbf{a}\phi_{\textrm{NN}}(\mathbf{x}),$ (1)
    where $\mathbf{1}\in\mathbb{R}^{r}$ is a vector of ones, and $\mathbf{a}\in\mathbb{R}^{r}$ is a vector of learnable weights.
Next for the decoder part, we project the latent state $\mathbf{z}$ back to the reconstructed system state following eq. 2,
$\hat{\mathbf{x}}=\psi_{\textrm{POD}}(\mathbf{z})(\mathbf{I}-\mathbf{B})+\psi_{\textrm{NN}}(\mathbf{z})\mathbf{B},$ (2)
where $\mathbf{I}$ is $Q\times Q$ identity matrix, $\mathbf{B}=\textrm{diag}(\mathbf{b})$, $\mathbf{b}\in\mathbb{R}^{Q}$, $\psi_{\textrm{POD}}$, and $\psi_{\textrm{NN}}$ are POD decoder and NN decoder, respectively. In addition to the parameters of NN encoder and decoder, both $\mathbf{a}\in\mathbb{R}^{r}$ and $\mathbf{b}\in\mathbb{R}^{Q}$ are trainable through gradient-based optimization as well. Hence, we name the above framework as learnable weighted hybrid autoencoder. It is important to note that such NN can be either fully-connected or convolutional.
[…]
To emphasize the role of learnable weight $\mathbf{a}$ and $\mathbf{B}$, we also implement a straightforward hybrid approach [22], which simply adds the latent states from POD and NN. Similarly, the state of the reconstructed system is the sum of the output of the POD decoder and the NN decoder. We name this hybrid approach as simple hybrid autoencoder. The difference of two approaches is further illustrated in fig. 1. As we shall see in the following sections, this comparison underscores the importance of using a weighted blend of the encoded and decoded spaces, as without it, the improvement over POD is at most incremental.
@@ 2.2 Koopman Forecasting
Koopman theory [23] states that any nonlinear dynamical system can be linearized by lifting into a (possible infinite-dimensional) space of observable functions. Hence, given a continuous dynamical system $\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})$, for any observable $\xi$, one defines the Koopman operator $\mathcal{K}$ as:
@@
where $\vec{\omega}$ refers to the frequency vector. With these assumptions the system state at any time $t$ is given by
@@
where $i$ indicates the timestep index and $\Delta t$ is the timestep size between consecutive snapshots. We utilize our proposed hybrid framework to learn the mapping from observable space to the physical state, while keeping their time evolution strategy unchanged, taking the same form as in eq. 2
$\psi_{\theta}(\Omega(\vec{\omega}t))=\psi_{\textrm{POD}}(\Omega(\vec{\omega}t))(\mathbf{I}-\mathbf{B})+\psi_{\textrm{NN}}(\Omega(\vec{\omega}t))\mathbf{B},$ (7)
@@ 2.3 Surrogate Modeling for Time-Dependent PDEs
Our surrogate modeling strategy involves the usage of the aforementioned techniques to obtain a latent representation and LSTM to determine its evolution over time. Training is divided into two stages. The dimensionality reduction framework is first trained to obtain a low-dimensional embedding of the system states, which subsequently serves as the training data for the time series prediction model to evolve the system in time. The advantage of training them separately is that it allows us to clearly demarcate the contribution of each stage to the overall surrogate model, which is particularly important as the crux of the current study is the introduction of a novel framework for dimensionality reduction.
@@ 3.1.1 Kuramoto-Sivashinsky (KS)
The governing equation of the KS equation is $\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}+\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{4}u}{\partial x^{4}}=0$, where $x\in[0,L_{x})$, $L_{x}=64\pi$, and periodic boundary conditions are assumed. The initial conditions comprise the sum of ten random sine and cosine waves $u(x,0)=\sum_{k=1}^{10}A_{k}\left(\sin\left(\frac{2\pi n_{k}x}{L_{x}}+\phi_{k}\right)+\cos\left(\frac{2\pi n_{k}x}{L_{x}}+\phi_{k}\right)\right)$, where $\forall k\in\{1,\ldots,10\},A_{k}\sim U(-1,1),\phi_{k}\sim U(0,2\pi)$ and wavenumber $n_{k}\sim U\{1,\ldots,6\}$, respectively. We employ Fourier decomposition in space and 2nd order Crank-Nicolson Adam-Bashforth semi-implicit finite difference scheme for temporal discretization. The data is generated for around 2000 time steps with a timestep of 0.01. Grid sizes of 512, 1024 and 2048 is used for the simulation. Some initial snapshots of the simulation that correspond to the transient phase are ignored.
For the NN encoder, we use a feedforward neural network with one hidden layer as shown in fig. 2, $2r$ number of neurons in the hidden layer. Hyperbolic tangent function is used for activation in all layers except the last linear layer. The NN decoder is symmetric to the encoder. After shuffling, the snapshot data is split into training and testing with a 7:3 ratio. Before training, we standardize the data by subtracting the mean and dividing by the standard deviation. All models are trained for 40k epochs with a batch size of 64.
Figure 2: NN-based encoder-decoder utilizing fully connected neural networks.
Figure 3: NN-based encoder-decoder utilizing convolutional neural networks.
@@ 3.1.2 Homogeneous isotropic turbulence (HIT)
We choose a deep convolutional autoencoder [26] as the NN part of the proposed framework, as shown in fig. 3. The details of the architecture pertaining to the NN based encoder and decoder and the training hyper parameters are shown in table 1. Without shuffling, we systematically sample the dataset by selecting every alternate snapshot to curate the training and testing data, each comprising 64 snapshots.
@@
Bottleneck Layer Fully Connected Linear Layer Output Layer Fully Connected Linear Layer
Activation Function Swish (except bottleneck layer) Activation Function Swish (except output layer)
Dropout 0.4 (All layers except bottleneck) Dropout 0.4 (All layers except output)
@@ 3.2.2 Flow Over Cylinder
We utilize the time evolution data of vorticity in a two-dimensional flow over a cylinder at $Re\approx 100$. The initial 50 temporal snapshots are used in training the Koopman-Decoder model, and the remaining 100 are kept out for testing. The details of the non-linear part of the decoder and training parameters are provided in table 3. The Koopman model is parametrized by two frequencies, essentially evolving the dynamics in a latent space of rank 2.

### v4

@@ Abstract
Representation learning for high-dimensional, complex physical systems aims to identify a low-dimensional intrinsic latent space, which is crucial for reduced-order modeling and modal analysis. To overcome the well-known Kolmogorov barrier, deep autoencoders (AEs) have been introduced in recent years, but they often suffer from poor convergence behavior as the rank of the latent space increases. To address this issue, we propose the learnable weighted hybrid autoencoder, a hybrid approach that combines the strengths of singular value decomposition (SVD) with deep autoencoders through a learnable weighted framework. We find that the introduction of learnable weighting parameters is essential — without them, the resulting model would either collapse into a standard POD or fail to exhibit the desired convergence behavior. Interestingly, we empirically find that our trained model has a sharpness thousands of times smaller compared to other models, which in turn enhances its robustness to input noise. Our experiments on classical chaotic PDE systems, including the 1D Kuramoto-Sivashinsky and forced isotropic turbulence datasets, demonstrate that our approach significantly improves generalization performance compared to several competing frameworks. Additionally, when combined with time series modeling techniques (e.g., Koopman operator, LSTM), the proposed technique offers significant improvements for surrogate modeling of high-dimensional multi-scale PDE systems.
@@ 1 Introduction
Computational fluid dynamics involves solving large dynamical systems with millions of degrees of freedom, resulting in significant computational overhead. In order to alleviate this problem, reduced-order modeling [1, 2] is widely used, which uses a smaller number of modes to provide an approximate solution at a lower computational expense. A crucial step in reduced-order modeling is the projection of high-dimensional system states to a reduced latent space [3]. The quality of projection can determine the overhead error for reduced-order modeling [4]. Linear dimensionality reduction techniques such as Proper Orthogonal Decomposition (POD) [5, 6, 7] are often used to create efficient representations of large-scale systems by projecting the solution manifold onto the space spanned by a set of linear orthonormal basis.
@@
Advances in deep learning techniques, such as deep autoencoders (AE) [8], capture intrinsic non-linear features for better compression, outperforming POD in low ranks, thus overcoming the well-known Kolmogorov barrier [9, 10]. Kolmogorov barrier is defined as a slow decay of the Kolmogorov $n$-width [11, 12], given by $d_{n}(\mathcal{M})=\inf_{\begin{subarray}{c}U_{n}\subset\mathcal{V},\dim(U_{n})=n\end{subarray}}\sup_{x(\cdot;t,\mu)\in\mathcal{M}}\inf_{\hat{x}\in U_{n}}\|x(\cdot;t,\mu)-\hat{x}\|$, where $\mathcal{M}$ is the solution manifold, $\mathcal{V}$ is the ambient Hilbert space, $U_{n}$ is any $n$-dimensional subspace of $\mathcal{V}$, $x(\cdot;t,\mu)$ is a solution to a parameterized PDE at time $t$ and parameter $\mu$, $\hat{x}$ is its projection onto $U_{n}$, and $\|\cdot\|$ denotes the norm in $\mathcal{V}$. The $n$-width quantifies the minimum worst-case projection error achievable by any $n$-dimensional linear subspace. This slow decay limits the best achievable error when using linear projection-based model reduction. Milano and Koumoutsakos [13] highlighted one of the first works to utilize a fully connected autoencoder to reconstruct the flow field, offering better performance compared to POD. Further studies have reported the usage of convolutional neural networks on 2D or 3D flow fields [14, 15, 16]. Such methods have been adopted in the fluid dynamics community to obtain a nonlinear model order reduction [17, 18], but they do not provide projection error convergence as the rank of the latent space increases. Recently, there have also been studies on using the hybrid approach [19, 20] combining POD with deep learning, by passing the latent space produced by POD to a neural network to find the corrections required to enhance reconstruction. These hybrid techniques have proven to enhance reconstruction beyond the capabilities of vanilla autoencoders. Unlike these approaches, which treat POD as a fixed preprocessing step and use neural networks only to adjust the POD output, in this work we propose a novel dimensionality reduction technique that combines traditional dimensionality reduction technique, i.e., POD, with deep learning techniques in a weighted manner at the encoder and decoder stage, where such weights of hybridization are also learned from data, to achieve a more effective dimensionality reduction. We also compare this approach with a straightforward hybrid technique, where a direct sum of POD and AE is utilized to construct the autoencoder, demonstrating the need of using learnable weighting parameters between POD and AE. Interestingly, our proposed approach obtains flat minima as opposed to other approaches, which contributes to the improved generalization and noise robustness.
@@
Building on our proposed framework, we also demonstrate the application of our framework to surrogate modeling tasks, particularly in the context of PDE system evolution. We explore two distinct applications of our framework. First, we integrate our framework with the Koopman operator, where our framework facilitates the mapping of Koopman embedding back to physical space, while learning a linear forward model to evolve the system states in a reduced space. Second, we leverage long-short-term memory (LSTM) for latent state evolution [21]. LSTMs are usually trained on the trajectories of reduced states, obtained via POD [22, 23] or a non-linear autoencoder [24, 25] to serve as a surrogate for the temporal evolution of latent states. The computational costs can be prohibitive when using LSTMs on full states, which can typically be in the range of millions for complex scenarios in fluid dynamics. By integrating our improved dimensionality reduction technique with a forward model for the latent states, we show that the overall framework leads to a more accurate prediction of the system dynamics and provide insights into the individual error contributions from the dimensionality reduction and time series model. This underscores the fact that the quality of the reduced representation is the primary bottleneck in achieving highly accurate predictions, as suboptimal dimensionality reduction inherently limit the effectiveness of any downstream models.
@@
The remainder of the manuscript is structured as follows. In Section 2 we present the details of our hybrid dimensionality reduction technique, Koopman operator used for long term forecast, and surrogate modeling for learning PDE dynamics with LSTM. Section 3 describes the datasets used in each of the tasks, network architecture, and training hyper-parameters. Section 4 provides a comprehensive evaluation of the proposed framework with respect to pure dimensionality reduction tasks and other downstream applications, comparing its performance against other techniques, and examining some key properties related to sharpness and noise robustness. Finally, Section 5 concludes the paper.
@@ 2.1 Dimensionality Reduction
    the neural network encoder with output dimension as $r$, denoted as $\phi_{\textrm{NN}}$. As shown in Equation 1, the latent state $\mathbf{z}\in\mathbb{R}^{r}$ is obtained by a weighted sum of POD projection and the output of encoder,
    $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{z}=(\mathbf{1}-\mathbf{a})\odot\phi_{\textrm{POD}}(\mathbf{x})+\mathbf{a}\odot\phi_{\textrm{NN}}(\mathbf{x}),$} (1)
    where $\mathbf{1}\in\mathbb{R}^{r}$ is a vector of ones, and $\mathbf{a}\in\mathbb{R}^{r}$ is a vector of learnable weights. $\phi$ is a function that produces a vector in $r$ dimensional subspace. The weight $\mathbf{a}$ is multiplied via a element-wise multiplication ($\odot$) to the latent representation produced by POD and NN.
Next, for the decoder part, we project the latent state $\mathbf{z}$ back to the reconstructed system state following Equation 2,
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{\mathbf{x}}=\psi_{\textrm{POD}}(\mathbf{z})\odot(\mathbf{1}-\mathbf{b})+\psi_{\textrm{NN}}(\mathbf{z})\odot\mathbf{b},$} (2)
where $\mathbf{1}\in\mathbb{R}^{Q}$ is a vector of ones, $\mathbf{b}\in\mathbb{R}^{Q}$, $\psi_{\textrm{POD}}$, and $\psi_{\textrm{NN}}$ are POD decoder and NN decoder, respectively. $\psi$ projects the latent vector into a physical space of dimension $\mathbb{R}^{N\times Q}$. Further an element wise multiplication ($\odot$) along the axis with dimension $Q$ multiplies the weights for the POD and NN part and combines them via an element wise addition. In addition to the parameters of NN encoder and decoder, both $\mathbf{a}\in\mathbb{R}^{r}$ and $\mathbf{b}\in\mathbb{R}^{Q}$ are trainable through gradient-based optimization as well. Hence, we name the above framework as learnable weighted hybrid autoencoder. It is important to note that such NN can be either fully-connected or convolutional.
Given the training dataset $\mathcal{D}=\{\mathbf{x}_{i}\}_{i=1}^{M}$, we trained the autoencoder by minimizing the mean-squared error (MSE) $\min_{\mathbf{\Theta},\mathbf{a},\mathbf{b}}\frac{1}{M}\sum_{i=1}^{M}\|\mathbf{x}_{i}-\hat{\mathbf{x}_{i}}\|^{2}$, where $\|\cdot\|$ is Frobenius norm, and $\mathbf{\Theta}$ refers to the set of the trainable parameters of neural network encoder $\phi_{\textrm{NN}}$ and decoder $\psi_{\textrm{NN}}$. $\mathbf{\Theta}$ is initialized using standard Kaiming initialization [26]. Motivated by Wang et al. [27], we choose to initialize $\mathbf{a}$ and $\mathbf{b}$ with zeros, leading to the proposed framework being equivalent to the classical POD at the beginning of neural network training. Thus, the model starts from the optimal linear encoder and becomes progressively nonlinear as the training proceeds. We choose Adam optimizer with learning rate of $10^{-4}$ for $\mathbf{\Theta}$ and $10^{-5}$ for $\mathbf{a}$ and $\mathbf{b}$.
To emphasize the role of learnable weight $\mathbf{a}$ and $\mathbf{b}$, we also implement a straightforward hybrid approach [28], which simply adds the latent states from POD and NN. Similarly, the state of the reconstructed system is the sum of the output of the POD decoder and the NN decoder. We name this hybrid approach as simple hybrid autoencoder. The difference of two approaches is further illustrated in Figure 1. As we shall see in the following sections, this comparison underscores the importance of using a weighted blend of the encoded and decoded spaces, as without it, the improvement over POD is at most incremental.
@@ 2.2 Koopman Forecasting
Koopman theory [29] states that any nonlinear dynamical system can be linearized by lifting into a (possiblly infinite-dimensional) space of observable functions. Hence, given a continuous dynamical system $\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})$, for any observable $\xi$, one defines the Koopman operator $\mathcal{K}$ as:
@@
where $\vec{\omega}$ denotes the frequency vector of dimension $N_{f}$, corresponding to the number of distinct frequencies considered. Consequently, the rank of the reduced representation, according to the formulation, becomes $r=2N_{f}$. With these assumptions the system state at any time $t$ is given by
@@
where $i$ indicates the timestep index and $\Delta t$ is the timestep size between consecutive snapshots. We utilize our proposed hybrid framework to learn the mapping from observable space to the physical state, while keeping their time evolution strategy unchanged, taking the same form as in Equation 2
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\psi_{\theta}(\Omega(\vec{\omega}t))=\psi_{\textrm{POD}}(\Omega(\vec{\omega}t))\odot(\mathbf{1}-\mathbf{b})+\psi_{\textrm{NN}}(\Omega(\vec{\omega}t))\odot\mathbf{b},$} (7)
@@ 2.3 Surrogate Modeling for Time-Dependent PDEs
Our surrogate modeling strategy involves the usage of the aforementioned techniques to obtain a latent representation and LSTM to determine its evolution over time. Training is divided into two stages. The dimensionality reduction framework is first trained to obtain a low-dimensional embedding of the system states, which subsequently serves as the training data for the time series prediction model to evolve the system in time[24, 31]. The advantage of training them separately is that it allows us to clearly demarcate the contribution of each stage to the overall surrogate model, which is particularly important as the crux of the current study is the introduction of a novel framework for dimensionality reduction.
@@ 3.1.1 Kuramoto-Sivashinsky (KS)
[…]
For the NN encoder, we use a feedforward neural network with one hidden layer with $2r$ number of neurons in the hidden layer. Hyperbolic tangent function is used for activation in all layers except the last linear layer. The NN decoder is symmetric to the encoder. After shuffling, the snapshot data is split into training and testing with a 7:3 ratio. Before training, we standardize the data by subtracting the mean and dividing by the standard deviation. All models are trained for 40k epochs with a batch size of 64.
@@ 3.1.2 Homogeneous isotropic turbulence (HIT)
We choose a deep convolutional autoencoder [33] as the NN part of the proposed framework. The details of the architecture pertaining to the NN based encoder and decoder and the training hyper parameters are shown in Table 1. Without shuffling, we systematically sample the dataset by selecting every alternate snapshot to curate the training and testing data, each comprising 64 snapshots.
@@
Bottleneck Layer Fully Connected Linear Output Layer Fully Connected Linear
Activation Function Swish (except bottleneck) Activation Function Swish (except output)
Dropout 0.4 (all except bottleneck) Dropout 0.4 (all except output)
@@ 3.2.2 Flow Over Cylinder
We utilize the time evolution data of vorticity in a two-dimensional flow over a cylinder at $Re\approx 100$[30]. The initial 50 temporal snapshots are used in training the Koopman-Decoder model, and the remaining 100 are kept out for testing. The details of the non-linear part of the decoder and training parameters are provided in Table 3. The Koopman model is parametrized by two frequencies, essentially evolving the dynamics in a latent space of rank 4.